# Some frameworks for dissipative evolution in multiscale non-equilibrium thermodynamics

### v1

@@ 3.2 GENERIC framework
Gradient dynamics generates the dissipative part of the GENERIC framework [31],[30], [60, 38, 59, 63],[6]
@@ 3.3 Metriplectic dynamics
can be seen as metriplectic dynamics (14). However, not all metriplectic systems can be reformulated by a dissipation potential [40, 35].
@@
3.4 Entropy production maximization
Other frameworks that are related to gradient dynamics are variational frameworks where the entropy production is maximized subject to some constraints. In the frameworks of Entropy Production Maximization [73, 74, 66, 56], the entropy production formula $J^{\alpha}X_{\alpha}$ is maximized subject to constraints given by a functional $\sigma(\mathbf{X})$. The functional actually represents the entropy production or a metric in the space of thermodynamic forces, similarly to the Steepest Entropy Ascent framework [3], where constraints are also imposed during the maximization.
@@ 3.6 A variational principle for GENERIC
The Onsager variational principle can be generalized to a variational principle for GENERIC [17, 18]. For state variables $\mathbf{x}\in M$, the action is given on the cotangent bundle $T^{*}M$ by
@@
where $\Xi$ is a convex dissipation potential, $\mathbf{x}$ is the position variable, $\mathbf{x}^{*}$ is the momentum variable, and $X_{H}$ is the Hamiltonian vector field $\{x,H\}$ for energy $H(\mathbf{x})$. The Euler-Lagrange equations are then
@@
where the first equation is the GENERIC equation (13).
However, to make the GENERIC equation closed within the state variables $\mathbf{x}$, the conjugate variables need to be identified with the derivatives of entropy, $\mathbf{x}^{*}=dS$, forming a section $dS(\mathbf{x})$ of the cotangent bundle $T^{*}M$. The right-hand side of Equations (20) represents the components of vector field $X_{\Psi}:T^{*}M\rightarrow TT^{*}M$ with $\Psi=-\Xi+\mathbf{x}^{*}\dot{\mathbf{x}}$,
@@ 4 Rayleigh dissipation potential
4.1 Extended GENERIC
$\left(\begin{array}[]{cc}\dot{\mathbf{x}}\\ \dot{\mathbf{w}}\end{array}\right)=\left(\begin{array}[]{cc}\mathbf{L}&K\\ -K^{T}&0\end{array}\right)\left(\begin{array}[]{cc}\mathbf{x}^{*}\\ \mathbf{w}^{*}\end{array}\right)+\left(\begin{array}[]{cc}0\\ \Xi_{\mathbf{w}^{*}}\end{array}\right)$ (26)
where $\mathbf{L}$ is the Poisson bivector expressing the Hamiltonian kinematics of $\mathbf{x}$ (i.e., $\{A,B\}=\langle A_{\mathbf{x}},\mathbf{L}B_{\mathbf{x}}\rangle)$, $K$ is a linear operator, $K^{T}$ is its transpose, $(\mathbf{x}^{*},\mathbf{w}^{*})$ are conjugate state variables, and $\Xi(\mathbf{w}^{*},\mathbf{x})$ is a dissipation potential. We assume that $\Xi$ is independent of $\mathbf{w}$ and, as a function of $\mathbf{w}^{*}$, it satisfies all the properties required of the dissipation potential. We directly verify that (26) is indeed a particular realization of the GENERIC equation, except for the Jacobi identity, which is not necessarily required in the presence of dissipation. If external forces are absent then there exists entropy $S(\mathbf{x},\mathbf{w})$ and $\dot{S}\geq 0$ provided $\mathbf{x}^{*}=S_{\mathbf{x}}$ and $\mathbf{w}^{*}=S_{\mathbf{w}}$.
We note that the second equation in (26) governs the time evolution of $\mathbf{w}^{*}$ provided we transform its left-hand side into $\dot{\mathbf{w}}^{*}$. We now assume that $\mathbf{w}$ evolves faster than $\mathbf{x}$; in the second equation in (26), we consider $\mathbf{x}$ as a fixed parameter. The second equation in (26) governing the fast time evolution of $\mathbf{w}^{*}$ becomes
$\dot{\mathbf{w}}^{*}=-\left(S_{\mathbf{w}\mathbf{w}}\right)^{-1}\mathfrak{R}_{\mathbf{w}^{*}}$ (27)
where
$\mathfrak{R}(\mathbf{w}^{*},\mathbf{x})=-\Xi(\mathbf{w}^{*},\mathbf{x})+\langle\mathbf{x}^{*},X\rangle$ (28)
is the rate thermodynamic potential called the Rayleigh potential; $X=K^{T}\mathbf{x}^{*}$. Consequently, $\dot{\mathfrak{R}}=-\langle\mathfrak{R}_{\mathbf{w}^{*}},\left(S_{\mathbf{w}\mathbf{w}}\right)^{-1}\mathfrak{R}_{\mathbf{w}^{*}}\rangle\geq 0$. The Hessian $S_{\mathbf{w}\mathbf{w}}$ is a negative definite operator. As for solutions of (27), we note that $\mathfrak{R}$ increases during the time evolution. As $t\rightarrow\infty$, solutions of (27) approach the maximum of $\mathfrak{R}$. The potential -$\mathfrak{R}$ plays the role of the Lyapunov function for the approach. The asymptotic solutions of (27) (i.e. solutions of $\mathfrak{R}_{\mathbf{w}^{*}}=0$) are
$\mathbf{w}^{*}=\mathfrak{R}^{{\dagger}}_{(\mathbf{w}^{*})^{{\dagger}}}((\mathbf{w}^{*})^{{\dagger}},\mathbf{x})|_{(\mathbf{w}^{*})^{{\dagger}}=\widehat{J}}$ (29)
With $\mathbf{w}^{*}$ given in (29), the equation in (26) that governs the time evolution of $\mathbf{x}$ becomes the GENERIC equation
$\dot{\mathbf{x}}=\mathbf{L}\mathbf{x}^{*}+K\mathfrak{R}^{{\dagger}}_{X}(X,\mathbf{x})|_{X=K^{T}\mathbf{x}^{*}}$ (30)
Indeed, we see that $\dot{S}(\mathbf{x})=\langle X,\mathfrak{R}_{X}\rangle\geq 0$. More detailed formulation of the extended GENERIC equation can be found in [36, 2].
@@ 4.3 Generalized formulation with Hamiltonian mechanics
which depends on the state variables $\mathbf{x}$ and energetic conjugates $\mathbf{x}^{\dagger}$ (to be identified with derivatives of energy $\mathbf{x}^{\dagger}=\frac{\delta E}{\delta\mathbf{x}}$). The density of the dissipation potential is denoted by $\mathfrak{r}$, and it can depend also on the spatial gradients of the conjugate fields. The time evolution is then given by
@@
For the energy conservation, it is required that $\mathfrak{r}$ does not explicitly depend on $s^{\dagger}$, as such a dependence could be used for instance to describe the radiative heat transfer with the environment. Energy is then conserved, as the evolution of the volumetric energy density $e(\mathbf{x})$ is of the divergence form,
@@
where $\mathbf{J}_{e}$ is the total energy flux generated by the Hamiltonian evolution (obtained from the evolution equations by the chain rule).
@@ 4.4 Gradient formulation
$\dot{\mathbf{q}}=$ $\frac{\delta\Xi}{\delta\mathbf{q}^{*}}=\frac{\partial\xi}{\partial\tilde{\mathbf{q}}^{*}}-\nabla\frac{\partial\xi}{\partial\widetilde{\nabla\mathbf{q}^{*}}}$, (43a)
@@
$\frac{\mathbf{q}^{\dagger}}{s^{\dagger}}\frac{\partial\xi}{\partial\tilde{\mathbf{q}}^{*}}+\frac{s^{*}}{s^{\dagger}}\frac{\partial\xi}{\partial\tilde{s}^{*}}=0$ (45)
is required for the energy conservation. This condition is analogical to condition (9) within gradient dynamics.
@@ 5.2.1 Hamiltonian picture
Near thermodynamic equilibrium, the dissipation potential it can be approximated by a quadratic functional,
@@ 5.2.2 Reduced dynamics
$R^{*}(H_{\mathbf{q}})=-H_{\mathbf{q}}\tilde{H}_{\mathbf{p}}(H_{\mathbf{p}})-R(\tilde{H}_{\mathbf{p}}(H_{\mathbf{q}})),$ (58)
@@ 6.2.1 Energetic Ehrenfest regularization
since $S_{x^{i}}L^{ki}=0$, so Casimirs are conserved. This behavior is suitable for instance in dissipative rigid-body dynamics, where the energy should decrease while keeping the Casimirs (the magnitude of angular momentum).
@@ 6.2.2 Entropic Ehrenfest regularization
so entropy, or any concave Casimir, is being produced (convex Casimirs are being reduced). This behavior can be used to generate dissipation in kinetic theory or fluid mechanics, where similar terms appear as in [69], where Laplacians of the state variables are added to all evolution equations in fluid mechanics.
6.3 Contact structure geometry

### v2

@@ 3.3 GENERIC framework
Gradient dynamics generates the dissipative part of the GENERIC framework [44, 43, 79, 52, 78, 82, 9]
@@ 3.4 Metriplectic dynamics
can be seen as metriplectic dynamics (16). However, not all metriplectic systems can be reformulated with a dissipation potential [54, 49].
@@
3.5 Entropy production maximization and Steepest Entropy Ascent
Other frameworks that are related to gradient dynamics are variational frameworks where the entropy production is maximized subject to some constraints. In the frameworks of Entropy Production Maximization [99, 100, 88, 74], the entropy production formula $J^{\alpha}X_{\alpha}$ is maximized subject to constraints given by a functional $\sigma(\mathbf{X})$. The functional actually represents the entropy production or a metric in the space of thermodynamic forces.
@@ 3.5 Entropy production maximization and Steepest Entropy Ascent
A relation between the entropy production maximization and the principle of large deviations with a quadratic rate function for metriplectic systems has been shown in [90].
A framework closely related to the entropy production maximization is the Steepest Entropy Ascent (SEA) [5], which has been shown to be essentially equivalent with the GENERIC framework with a quadratic dissipation potential [69], that is with metriplectic systems. SEA represents a very general framework for dissipative evolution with many applications [6].
In SEA, a strongly nondegenerate metric tensor $\mathbf{g}$ on the state space $M$ is introduced, which provides a scalar product $(\mathbf{X},\mathbf{Y})=\mathbf{g}(\mathbf{X},\mathbf{Y})$ of two vectors $\mathbf{X}$ and $\mathbf{Y}$ in the tangent space of the state space. Then, in the presence of constraints $c_{\alpha}(\mathbf{x})$, the entropy gradient is first projected onto the subspace orthogonal to the gradients of the constraints,
$ds^{c}=ds-\sum_{\alpha}\lambda^{\alpha}dc_{\alpha},$ (18)
where the Lagrange multipliers $\lambda^{\alpha}$ are determined by the orthogonality condition $(ds^{c},dc_{\alpha})=0$.
The SEA evolution is then given by the sum of the reversible vector field $\mathbf{X}_{H}$ and irreversible vector field $\mathbf{Y}_{S}$ where the latter is given by the steepest ascent of the entropy, $\mathbf{Y}_{S}=\frac{1}{\tau}\mathbf{g}^{-1}ds^{c}$, where $\tau$ is a positive dimensional prefactor. The irreversible evolution then automatically respects the constraints, since $dc_{\alpha}(\mathbf{Y}_{S})=(ds^{c},dc_{\alpha})=0$. Moreover, the entropy production is non-negative, $\dot{S}=ds(\mathbf{Y}_{S})=(ds^{c},ds^{c})/\tau\geq 0$.
The SEA evolution can be also reformulated as a variational principle, where the entropy production is maximized subject to the constraint that the magnitude of the irreversible vector field $\mathbf{Y}_{S}$ is fixed,
$\max_{\mathbf{Y}_{S}}ds^{c}(\mathbf{Y}_{S})\quad\text{subject to}\quad\mathbf{g}(\mathbf{Y}_{S},\mathbf{Y}_{S})=\text{const},$ (19)
where the dimensional prefactor $\tau$ then plays the role of the Lagrange multiplier of the constraint. In other words, the SEA framework can be seen as quadratic entropy production maximization that automatically respects the constraints. Alternatively, if the reversible vector field $\mathbf{X}_{H}$ is Hamiltonian and does not affect the total entropy, the SEA framework can be seen as a metriplectic system (or GENERIC with quadratic dissipation potential).
Therefore, SEA has the same relation with GENERIC in its gradient form (15) as metriplectic systems (16). Since the whole metric tensor $\mathbf{g}$ must be supplied in SEA, it makes it more general than GENERIC in the gradient form, as not all metric tensors are Hessians of dissipation potentials. However, instead of one non-linear functional $\Xi$ (dissipation potential), SEA requires $n\cdot(n+1)/2$ functions (components of the metric tensor) to be supplied, where $n$ is the number of state variables. Moreover, the principle large deviations [66, 80, 70] leads to gradient dynamics with a generally non-quadratic dissipation potential.
@@ 3.7 A variational principle for GENERIC
There are several variational formulations of the GENERIC framework. One possibility is to use the formulation within contact geometry [46]. Another option is the Onsager variational principle for GENERIC [26, 27]. In the case of quadratic dissipation potential, also SEA provides a variational formulation of GENERIC [69]. Finally, the principle of large deviations leads to a variational formulation of GENERIC [66, 80, 70, 59].
Here, we present another variational formulation of GENERIC which is formulated for paths on the cotangent bundle $T^{*}M$ of the state space $M$. Assume that the Hamiltonian vector field $X_{H}:M\rightarrow TM$ is given by the Poisson bracket and energy, $X_{H}=\{\mathbf{x},H\}$. The cotangent bundle $T^{*}M$ is then parametrized by the state variables $\mathbf{x}$ and their conjugate variables $\mathbf{x}^{*}$. Then, stationary points of action
@@
where $\Xi$ is a convex dissipation potential are given by Euler-Lagrange equations
@@
where the first equation is the GENERIC equation (15) without the identification of $\mathbf{x}^{*}$ with the gradient of entropy.
To make the GENERIC equation closed within the state variables $\mathbf{x}$, the conjugate variables need to be identified with the derivatives of entropy, $\mathbf{x}^{*}=dS$, forming a section $dS(\mathbf{x})$ of the cotangent bundle $T^{*}M$. The right-hand side of Equations (24) represents the components of vector field $X_{\Psi}:T^{*}M\rightarrow TT^{*}M$ with $\Psi=-\Xi+\mathbf{x}^{*}\dot{\mathbf{x}}$,
@@
Note that the integrand in action (23) actually depends on $\dot{\mathbf{x}}$, but it can be rewritten as $\Xi(\mathbf{x},\mathbf{x}^{*})dt+\mathbf{x}^{*}d\mathbf{x}-\mathbf{x}^{*}X_{H}dt$, so we do not need to explicitly include the tangent direction $\dot{\mathbf{x}}$ into the underlying geometry. Alternatively, we could go fully to the contact geometry [46].
@@ 4.2 Generalized formulation with Hamiltonian mechanics
which depends on the state variables $\mathbf{x}$ and energetic conjugates $\mathbf{x}^{\dagger}$ (to be identified with derivatives of energy $\mathbf{x}^{\dagger}=\frac{\delta E}{\delta\mathbf{x}}$), see [24]. The density of the dissipation potential is denoted by $\mathfrak{r}$, and it can depend also on the spatial gradients of the conjugate fields. The time evolution is then given by
@@
For energy conservation, it is required that $\mathfrak{r}$ does not explicitly depend on $s^{\dagger}$, as such a dependence could be used for instance to describe the radiative heat transfer with the environment. Energy is then conserved, as the evolution of the volumetric energy density $e(\mathbf{x})$ is of the divergence form,
@@
where $\mathbf{J}_{e}$ is the total energy flux generated by the Hamiltonian evolution (obtained from the evolution equations by the chain rule). Note, however, that in order to guarantee the chain rule, all fields and potentials must be sufficiently regular [29, 91].
@@ 4.3 Gradient formulation
$\dot{\mathbf{q}}=$ $\frac{\delta\Xi}{\delta\mathbf{q}^{*}}=\frac{\partial\xi}{\partial\tilde{\mathbf{q}}^{*}}-\nabla\cdot\frac{\partial\xi}{\partial\widetilde{\nabla\mathbf{q}^{*}}}$, (42a)
@@
$\frac{\mathbf{q}^{\dagger}}{s^{\dagger}}\cdot\frac{\partial\xi}{\partial\tilde{\mathbf{q}}^{*}}+\frac{s^{*}}{s^{\dagger}}\frac{\partial\xi}{\partial\tilde{s}^{*}}=0$ (44)
is required for the energy conservation. This condition is analogous to condition (8) within gradient dynamics.
@@ 4.4 Single generator formalism
4.5 Extended GENERIC
An alternative route towards the Rayleigh dissipation potential is to start from the GENERIC equation and to extend it by adding additional state variables $\mathbf{w}$ that evolve faster than the original state variables $\mathbf{x}$. The fast evolution of $\mathbf{w}$ then leads to the Rayleigh dissipation potential for the slow evolution of $\mathbf{x}$.
$\left(\begin{array}[]{cc}\dot{\mathbf{x}}\\ \dot{\mathbf{w}}\end{array}\right)=\left(\begin{array}[]{cc}\mathbf{L}&\mathcal{K}^{\dagger}\\ -\mathcal{K}&0\end{array}\right)\left(\begin{array}[]{cc}\mathbf{x}^{*}\\ \mathbf{w}^{*}\end{array}\right)-\left(\begin{array}[]{cc}0\\ \Xi_{\mathbf{w}^{*}}\end{array}\right)$ (48)
where $\mathbf{L}$ is the Poisson bivector expressing the Hamiltonian kinematics of $\mathbf{x}$ (i.e., $\{A,B\}=\langle A_{\mathbf{x}},\mathbf{L}B_{\mathbf{x}}\rangle)$, $\mathcal{K}$ is a linear operator, $\mathcal{K}^{\dagger}$ is its adjoint, $(\mathbf{x}^{*},\mathbf{w}^{*})$ are conjugate state variables, and $\Xi(x,X,\mathbf{w}^{*})$ is a dissipation potential. We assume that $\Xi$ is independent of $\mathbf{w}$ and, as a function of $\mathbf{w}^{*}$, it satisfies all the properties required of the dissipation potential [82]. Equation (48) is indeed a particular realization of the GENERIC equation, except for the Jacobi identity, which is not necessarily required in the presence of dissipation. If external forces are absent, then there exists a thermodynamic potential $\Phi(\mathbf{x},\mathbf{w})$ and $\dot{\Phi}\leq 0$ provided $\mathbf{x}^{*}=\Phi_{\mathbf{x}}$ and $\mathbf{w}^{*}=\Phi_{\mathbf{w}}$, typically $\Phi(\mathbf{x})=-S(\mathbf{x},\mathbf{w})+\frac{1}{T_{0}}E(\mathbf{x},\mathbf{w})$ where $T_{0}$ is a fixed parameter (the equilibrium temperature).
We note that the second equation in (48) governs the time evolution of $\mathbf{w}^{*}$ provided we transform its left-hand side into $\dot{\mathbf{w}}^{*}$. We now assume that $\mathbf{w}$ evolves faster than $\mathbf{x}$, so in the second equation in (48) $\mathbf{x}$ is treated as a fixed parameter for the moment. The second equation in (48) governing the fast time evolution of $\mathbf{w}^{*}$ becomes
$\dot{\mathbf{w}}^{*}=-\Phi_{\mathbf{w}\mathbf{w}}\mathfrak{R}_{\mathbf{w}^{*}}$ (49)
where
$\mathfrak{R}(x,\mathbf{w}^{*},X)=\Xi(\mathbf{w}^{*},X)+\langle\mathcal{K}\mathbf{x}^{*},\mathbf{w}^{*}\rangle$ (50)
is the rate thermodynamic potential called the Rayleigh potential and $X=\mathcal{K}\mathbf{x}^{*}$ a thermodynamic force. Consequently, $\dot{\mathfrak{R}}=-\langle\mathfrak{R}_{\mathbf{w}^{*}},\Phi_{\mathbf{w}\mathbf{w}}\mathfrak{R}_{\mathbf{w}^{*}}\rangle\leq 0$, as the Hessian $\Phi_{\mathbf{w}\mathbf{w}}$ is a positive definite operator, assuming sufficient regularity and locality. As for solutions of (49), we note that $\mathfrak{R}$ decreases with $t\rightarrow\infty$ and potential $\mathfrak{R}$ then plays the role of the Lyapunov function (or S-function).
When we construct the dual dissipation potential $\Xi^{*}(\mathbf{x},X,J_{\mathbf{w}})$ by the Legendre transformation $\Xi^{*}=-\Xi+\langle J_{\mathbf{w}},\mathbf{w}^{*}\rangle$, $J_{\mathbf{w}}=\Xi_{\mathbf{w}^{*}}$, $\mathbf{w}^{*}=\Xi^{*}_{J_{\mathbf{w}}}$, the quasi-equilibrium solution of (48) becomes $\mathbf{w}^{*}=\Xi^{*}_{J_{\mathbf{w}}}\Big|_{J_{\mathbf{w}}=-X}$, and the remaining equation for the slow evolution of $\mathbf{x}$ becomes
$\dot{\mathbf{x}}=\mathbf{L}\mathbf{x}^{*}+\mathcal{K}^{{\dagger}}\Xi^{*}_{J_{\mathbf{w}}}\Big|_{J_{\mathbf{w}}=-X},$ (51)
which has the GENERIC form for the $\mathbf{x}$ variable. In particular, we see that $\dot{\Phi}(\mathbf{x})=-\langle X,\Xi^{*}_{X}\rangle\leq 0$. More detailed formulation of the extended GENERIC equation can be found in [50, 3].
@@ 5.2.1 Hamiltonian picture
Note, however, that the Lagrangian may be singular, and thus the Legendre transformation does not need to be diffeomorphic. In that case, one may use the framework of Tulczyjew triples [95], using Lagrangian submanifolds. Other possibility is the use of the slope transformation [21, 22, 56], where the Legendre-transformed potentials become graphs (or multivalued functions).
Near thermodynamic equilibrium, the dissipation potential can be approximated by a quadratic functional,
@@ 5.2.2 Reduced dynamics
$R^{*}(H_{\mathbf{q}})=-H_{\mathbf{q}}\tilde{H}_{\mathbf{p}}(H_{\mathbf{q}})-R(\tilde{H}_{\mathbf{p}}(H_{\mathbf{q}})),$ (61)
@@ 6.2.1 Energetic Ehrenfest regularization
since $S_{x^{i}}L^{ki}=0$, so Casimirs are conserved. This behavior is suitable for instance in dissipative rigid-body dynamics,
$\dot{\mathbf{m}}=\mathbf{m}\times\boldsymbol{\omega}+\frac{\tau}{2}\mathbf{m}\times\left(\mathbb{I}^{-1}(\mathbf{m}\times\boldsymbol{\omega})\right)$ (74)
where $\mathbf{m}$ is the angular momentum in a reference frame fixed with the body, $\mathbb{I}^{-1}=d^{2}H$ is the inverse of the inertia tensor, and $\boldsymbol{\omega}=\mathbb{I}^{-1}\mathbf{m}$ is the angular velocity. Energy then decreases in time while keeping the Casimirs (the magnitude of angular momentum).
@@ 6.2.2 Entropic Ehrenfest regularization
so entropy, or any concave Casimir, is being produced (convex Casimirs are being reduced). This behavior can be used to generate dissipation in kinetic theory,
$\partial_{t}f=-\frac{p_{i}}{m}\frac{\partial f}{\partial r^{i}}+\frac{\tau}{2}\frac{p_{k}}{m}\frac{p_{l}}{m}\frac{\partial^{2}f}{\partial r^{k}\partial r^{l}},$ (78)
or fluid mechanics, where similar terms appear as in [92], where Laplacians of the state variables are added to all evolution equations in fluid mechanics.
7 Comparison of the Port-Hamiltonian and the Extended GENERIC views of Thermodynamics of Externally Driven Systems
When carrying thermodynamics from externally unforced to externally driven systems the state variables shift from the base space to its tangent and cotangent spaces. This result has appeared in the context of the GENERIC formulation of thermodynamics, and it has also appeared in its Port-Hamiltonian (PH) formulation [96]. In the former it is because thermodynamics is associated with dynamics describing the approach to fixed points. In externally unforced systems, the fixed points are thermodynamic equilibrium states in the base space. In externally driven systems, the fixed points are vector fields generating the time evolution in reduced dynamical theories, involving less detail. From the PH viewpoint, the necessity to turn attention to tangent and cotangent spaces has appeared in the attempt to bring the power of thermodynamics to the control theory [65, 97].
We hope that the comparison of the Extended GENERIC (48) with the PH thermodynamics that we formulate below will help to combine and enlarge the pools of insights that emerged independently in both GENERIC and PH views of thermodynamics.
Vector space $W$ with elements $w=(\xi,w_{R},w_{ext})\in W$ plays the role of state variables in PH-thermodynamics. The first vector is $\xi\in T_{x}M$, where $M$ is the state space GENERIC-thermodynamics; $x\in M$. The vector $w_{R}$ is called a resistive state variable and $x_{ext}$ is a vector expressing an external influence (like control and boundary). The vector space $W^{*}$ with elements $w^{*}=(\xi^{*},w^{*}_{B},w^{*}_{ext})\in W^{*}$ is dual to $W$. The element $\xi^{*}$ is physically interpreted as gradient of the energy $H(x)$, i.e. $\xi^{*}=H_{x}$. To each pair $(w,w^{*})$, we associate $\langle w^{*},w\rangle\in\mathbb{R}$.
Arguments developed in the control theory and thermodynamics [65] led to the following result. Processes that are compatible with the Hamiltonian dynamics and thermodynamics are restricted to a submanifold $\mathcal{D}\subset W\oplus W^{*}$ determined by (see e.g. Eq.(30) in ([97]))
$\left(\begin{array}[]{ccc}\dot{x}\\ w_{R}\\ w_{ext}\end{array}\right)=\left(\begin{array}[]{ccc}\mathbb{L}&\mathcal{K}_{R}(x)&\mathcal{K}_{ext}(x)\\ -\mathcal{K}_{R}^{\dagger}(x)&0&0\\ -\mathcal{K}_{ext}^{\dagger}(x)&0&0\end{array}\right)\left(\begin{array}[]{ccc}H_{x}\\ w_{R}^{*}\\ w_{ext}^{*}\end{array}\right)$ (79)
and
$R(w_{R},w^{*}_{R})=0\,\,\text{guaranteeing}\,\,\langle w^{*}_{R},w_{R}\rangle\leq 0\,\forall\,\,(w^{*}_{R},w_{R})$ (80)
where $\mathcal{K}_{R}(x)$ and $\mathcal{K}_{ext}(x)$ are linear operators, $R$ is in general a nonlinear function of $(w_{R},w^{*}_{R})$.
Now we compare (79) with (48). With $(w_{R},w_{R}^{*},w_{ext},w_{ext}^{*})$ absent in (79) and $w$ absent in (48), the first equations in both (79) and (48) are the same. To compare the second equations in (79) and (48) we interpret $w_{R}$ as an element $\dot{\zeta}_{R}$ of the tangent space $TN_{R}$ of the space $N_{R}$ with elements $\zeta_{R}\in N_{R}$. With $(w_{ext},w_{ext}^{*})$ absent, the first two equations in (79), describing now the time evolution in an extended state space $M\times N_{R}$, become
$\left(\begin{array}[]{cc}\dot{x}\\ \dot{\zeta}_{R}\end{array}\right)=\left(\begin{array}[]{cc}\mathbb{L}&\mathcal{K}_{R}(x)\\ -\mathcal{K}_{R}^{\dagger}(x)&0\end{array}\right)\left(\begin{array}[]{cc}H_{x}\\ w_{R}^{*}\end{array}\right).$ (81)
This reinterpretation of $w_{R}$ allows us to express the resistivity in a way that is physically more transparent than the requirement (80).
We assume that the resistivity is generated in the time evolution of $\zeta_{R}$. This assumption is expressed mathematically by modifying the second equation in (81). We add to its right-hand side the dissipative term in the same way as we do in (48), and the modified second equation in (81) becomes
$\dot{\zeta}_{R}=-\mathcal{K}^{T}_{R}H_{x}+\Xi_{w^{*}}(w^{*},x).$ (82)
If the dissipation is strong (i.e. $\Xi_{w^{*}}(w^{*},x)$ is dominant on the right-hand side of (82)), then $\zeta_{R}$ evolves rapidly to
$\dot{\zeta}_{R}=0=-K^{T}_{R}H_{x}+\Xi_{w^{*}}(w^{*},x).$ (83)
Summing up, we have replaced the two equations
$w_{R}$ $=$ $-K_{R}^{T}(x)H_{x}(x)$
$R(w_{R},w^{*}_{R})$ $=$ $0$ (84)
that allow to eliminate the resistive state variables $(w_{R},w^{*}_{R})$ from the PH equations (79) with Eq.(83).
If in the third equation in (79) we interpret $w_{ext}$ as an element $\dot{\zeta}_{ext}$ of the tangent space $TN_{ext}$ of the space $N_{ext}$ with elements $\zeta_{ext}\in N_{ext}$ then the PH formulation (79), (80) of thermodynamics of driven systems becomes completely equivalent to the Extended GENERIC formulation (48) describing the time evolution in the state space $M\times N_{R}\times N_{ext}$.
In both the formulations of extensions of thermodynamics from externally unforced to externally driven systems, the extended state variables are rate state variables (i.e. elements of tangent and cotangent bundles). In the Extended GENERIC formulation they emerge as elements of tangent and cotangent spaces of an extended state space. From the physical point of view, the enlargement of the base space means that the external influences are internalized. In the PH formulation the state space (the base space of the bundles) remains unchanged, only its tangent and cotangent spaces are enlarged.
This type of extension has also been investigated in the GENERIC formulation in [1] in the context of chemical kinetics. From the mathematical point of view, the extension in [1] involves the structure of Lie algebroids [86], in the PH extension it involves the Dirac structure [17] (the submanifold $\mathcal{D}\subset W\oplus W^{*}$ specified explicitly in (79) appears to be a Dirac structure on $W\oplus W^{*}$). We intend to explore both the physical and the mathematical aspects of these two views of extensions in the future.
8 Contact structure geometry